OpenStudy (anonymous):

Please Someone Help... The equation below shows the relationship between the temperature in degrees Celsius, C, and degrees Fahrenheit, F. F = 9 Over 5 C + 32 Which of the following formulas correctly solves for C? Options: a. C = 5F - b. C = - 32 c. C = (F - 32) d. C = (F - 32)

OpenStudy (anonymous):

@SkiTTleoooo47

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

Nope Im Confused.

OpenStudy (skullpatrol):

Use the idea of "balance" to become unconfused.

OpenStudy (skullpatrol):

Whatever you do to one side of the equation, you must do the same to the other, get the idea? Like a weight balance.

OpenStudy (anonymous):

Okay.

OpenStudy (skullpatrol):

$$\Huge F = \frac{9}5 C + 32$$ $$\Huge F-32 = \frac{9}5 C + 32-32$$ $$\Huge F-32 = \frac{9}5 C $$

OpenStudy (anonymous):

Can You Explain ?

OpenStudy (skullpatrol):

I subtracted 32 from BOTH sides.

OpenStudy (skullpatrol):

Now I will multiply BOTH sides by what?

OpenStudy (skullpatrol):

$$\Huge \frac{5}9\cdot(F-32) =\frac{5}9 \cdot\frac{9}5 C $$ $$\Huge \frac{5}9\cdot(F-32) = C $$

OpenStudy (anonymous):

Thanks

OpenStudy (skullpatrol):

Thanks for asking :)